# COMET: Co-Optimization of a CNN Model using Efficient-Hardware OBC Techniques

### v2

@@ III-D1 Hardware and Time Complexities
The theoretical estimates of hardware complexity (in terms of adders and 2-to-1 multiplexers) and time complexity (in terms of CPD) for the proposed LUT architectures are listed in Table I. These are estimated with the decomposition $K=p\times q$, to ensure fair comparison across designs. The Parallel LUT incurs the highest hardware cost due to exponential growth in both adders and multiplexers with respect to $q$. The Shared LUT reduces the adder count by reusing partial computations but retains similar CPD. The Split LUT lowers adder usage further, though this comes with an increase in multiplexers. Compared to the shared LUT, the Hybrid LUT offers the best balance, achieving linear adder scaling and reduced multiplexer count with respect to $q$. Notably, its CPD also remains low due to shallow adder trees. For a consistent comparison between different proposed LUT techniques, we set $q=4$, which implies $K=4p$. To validate these findings, all LUT architectures are implemented on FPGA with 8-bits symmetric operands for evaluation, as discussed in the following subsection.
@@ III-D2 LUT and FF Usage vs Clock Frequency
Fig. 7 shows the resource utilization of the proposed LUT architectures implemented on FPGA across a clock frequency range of 100–800 MHz in 100 MHz steps. This evaluation examines how each design behaves as the clock frequency increases. LUT usage remains largely stable when timing constraints are easily met. At higher frequencies, however, FPGA synthesis and place-and-route tools introduce additional optimization logic—such as logic duplication and register insertion—to satisfy tighter timing requirements. This increase reflects tool-driven overhead rather than the intrinsic LUT requirements of each design. To ensure consistent comparisons, the frequency range is restricted. All designs exhibit LUT usage consistent with the complexities listed in Table I. The Parallel LUT shows the highest LUT usage due to its large number of adders and multiplexers. The Shared LUT reduces the LUT count by reusing partial computations, achieving moderate savings over Parallel LUT. The Split LUT reduces adders but raises logic utilization from added 2’s complementers; for $K=32$, performance is slightly worse due to the extra 2’s complement operations. The Hybrid LUT demonstrates the most efficient resource usage overall owing to optimized adders and minimal multiplexer overhead. Small fluctuations in LUT usage may occur across all designs due to synthesis heuristics and routing decisions, but these variations are minor.
Unlike LUT usage, flip-flop (FF) utilization remains largely consistent across all LUT designs, as shown in Fig. 8. This is because LUTs are primarily combinational, while FFs are used to store input/output feature maps and intermediate pipeline values, which are independent of the LUT architecture. When the kernel size $K$ increases, the datapath widens, requiring more FFs to hold intermediate results. This increase is similar across all designs, as $K$ affects sequential storage rather than the combinational structure of the LUTs.
@@ III-D3 Power Consumption vs Clock Frequency
Accurate power estimation requires realistic activity data, as post-synthesis estimates often miss switching and routing effects in complex LUTs. In Xilinx Vivado, the process is: (1) run functional simulation; (2) generate a .saif file with signal toggles; (3) import it into power analysis; (4) combine toggles with utilization, timing, and clock frequency. This yields power figures that capture actual activity and frequency behavior. Using this procedure, power consumption increases with clock frequency for all LUT architectures due to higher switching activity and dynamic power. The slope of the power versus frequency curve varies across designs, reflecting differences in logic complexity, routing, and the number of active switching elements. The Parallel LUT exhibits the steepest slope because its large number of adders and multiplexers produces more toggling per cycle. The Shared LUT shows low power for $K=8$ because extensive computation reuse reduces switching activity, despite slightly higher combinational complexity. The Hybrid LUT, while consuming slightly more power than the Shared LUT at small $K$ due to active logic per patch, achieves more consistent power for $K=16$ and the lowest power at $K=32$ by efficiently pairing adders and minimizing multiplexers, thereby reducing toggling per operation. The Split LUT has shallower slopes than Parallel LUTs, as reduced adder counts lower overall switching activity. These trends show that power depends on both logic complexity and switching/data reuse.
TABLE II: Description of the Original and Modified LeNet‑5
Layer Original [34] Modified
Conv1 $5\times 5$ conv, 6 ch, tanh $5\times 5$ conv, 6 ch, ReLU
Pool1 $2\times 2$ avg pool, stride 2, tanh $3\times 3$ conv, stride 2, ReLU
Conv2 $5\times 5$ conv, 16 ch, tanh $5\times 5$ conv, 16 ch, ReLU
Pool2 $2\times 2$ avg pool, stride 2, tanh $3\times 3$ conv, stride 2, ReLU
GAP global avg pool on 16 ch
FC1 dense $400\!\to\!120$, tanh dense $16\!\to\!32$, ReLU
FC2 dense $120\!\to\!84$, tanh dense $32\!\to\!10$, softmax
FC3 dense $84\!\to\!10$, softmax
@@
IV Model Considerations, Accuracy
and Evaluation of the Proposed OBC-GEMM
IV-A Model Considerations
To enable efficient FPGA implementation, we apply some lightweight modifications to the baseline LeNet-5 architecture [34], as listed in Table II:
  • The original use of tanh activations introduces saturation and requires scaling. Replacing them with ReLU not only yields higher sparsity but also enables simpler fixed-point implementation.
  • The original $2\times 2$ average-pooling layers neither learn parameters nor map cleanly onto the GEMM datapath. Replacing them with learnable $3\times 3$ strided convolutions unifies computation [35], increases arithmetic intensity, and better exposes the advantages of the proposed design.
  • LeNet-5’s two large FC layers dominate the parameter count. Substituting them with a GAP layer followed by a compact $32$-unit FC drastically reduces memory traffic while preserving accuracy [21].
  • One-side zero padding is inserted before each strided-convolution down-sampling step to keep spatial dimensions consistent, simplifying address generation and buffer reuse.
These modifications are adopted to demonstrate that the OBC-GEMM core applies cleanly to modern workloads. With fixed quantization, frequency, and tiling, they remain orthogonal to throughput. Specifically, replacing $2\times 2$ average pooling with $3\times 3$ stride-2 convolutions increases MACs per inference, while swapping tanh for ReLU and adding zero padding incur negligible MAC costs. GAP reduces parameters and memory traffic but not MACs per cycle. As in [20], these modifications mainly affect accuracy and memory; therefore, it is necessary to explore their impact in fixed-point implementations.
Figure 9: Accuracy of original and modified LeNet-5 under QAT with two sets of input bit-widths ($B_{1}$) and varying weight bit-widths ($B_{2}$).
IV-B Accuracy Comparison of LeNet-5 vs Modified LeNet-5
To evaluate the impact of lightweight architectural modifications under quantization-aware training (QAT), we compared the baseline LeNet-5 with a modified variant on MNIST, using padded $32\times 32$ images. Both models were trained with fixed-point arithmetic, fixing input bitwidth at $B_{1}=16$ or $B_{1}=8$, while varying weight bitwidth $B_{2}$. Results are summarized in Fig. 9.
For $B_{1}=16$ (top subplot), the baseline LeNet-5 achieved $91.96\%$$92.58\%$, with no consistent improvement as $B_{2}$ increased. In contrast, the modified network consistently outperformed the baseline, reaching $98.28\%$ at $B_{2}=4$ and $99.16\%$ at $B_{2}=16$. Gains were largest at mid-range bitwidths, where accuracy was nearly seven points higher. This shows that under higher input bitwidth, the modified architecture is far more resilient to weight quantization, sustaining near–floating-point performance even at low bit-widths.
With reduced input bitwidth ($B_{1}=8$, bottom subplot), the baseline degraded sharply at low $B_{2}$, dropping to $83.18\%$ at 2 bits and recovering to $94.34\%$ at 8 bits. The modified model was even more sensitive at $B_{2}=2$ ($64.32\%$), but rebounded strongly thereafter, achieving $98.22\%$ at $B_{2}=4$ and $98.60\%$ at $B_{2}=8$, exceeding the baseline by over four points at higher bitwidths.
Overall, two key insights emerge. First, replacing tanh with ReLU and introducing stride-2 convolutions significantly improve quantization robustness, particularly under moderate-to-high bitwidth. Second, while the modified LeNet-5 is vulnerable at ultra-low bitwidth ($B_{2}=2$ with $B_{1}=8$), performance recovers quickly with modest bit-width increases. These results confirm that careful architectural adjustments enable efficient hardware deployment, where reduced bitwidth yields memory savings, lower power use, and faster inference.
IV-C Evaluation of the OBC-GEMM Module
Figure 10: LUTs/$F$s usage and power of OBC-GEMM for different $(B_{1},B_{2})_{S}$ with $S\in\{A,B,AB\}$ indicates Scheme A or B or Symmetric case - AB values with hybrid and split LUT techniques at 50 MHz and 200 MHz.
For evaluation purposes, we discuss the FPGA implementation results corresponding to the accuracy outcomes for input bitwidth $B_{1}=16$ while varying the weight bitwidth $B_{2}$; however, the analysis could be further expanded to include $B_{1}=8$. The proposed OBC-GEMM module naturally supports operands with asymmetric quantization under Scheme A and Scheme B, as described in Section II. In this study, we fix the input bit-width at $B_{1}=16$ and vary the weight bit-width $B_{2}$ according to the accuracy results shown in Fig. 9. The case of $B_{2}=4$ with $B_{1}=16$ under Scheme A is omitted due to significant accuracy degradation, but the corresponding $(16,4)_{B}$ configuration is retained to highlight Scheme B’s behavior when weights are aggressively quantized.
Under asymmetric quantization, both proposed schemes follow $B_{1}>B_{2}$. In Scheme A, the SA unit operates for $B_{1}$ clock cycles corresponding to the input bits, while the partial products of the weights are generated using efficient hardware techniques, reducing arithmetic complexity and overall hardware resource usage. In contrast, Scheme B (Fig. 5) bit-slices the weights with bit-width $B_{2}$ through a PISO for address generation, while inputs in OBC form are used to generate the LUT data. Here, the data remain in the SA unit for $B_{2}$ clock cycles, enabling faster output, but the wider LUT data ($B_{1}$ bits) increase logic utilization and power consumption. The $(16,4)_{B}$ case illustrates this trade-off in its most extreme form: very short SA operation but significantly wider LUT paths. After integrating the Split and Hybrid LUTs and implementing both schemes, the results are summarized in Fig. 10. To evaluate the effect of timing constraints, we vary $B_{2}$, and $K$ at two operating frequencies: 50 MHz and 200 MHz. At 50 MHz, the designs meet timing with minimal replication, while at 200 MHz additional logic and routing are required to satisfy stricter constraints. Under both conditions, Scheme A and Scheme B exhibit consistent resource–performance trade-offs, and the choice of $(B_{1},B_{2})$ directly governs efficiency. The detailed effects on LUT usage, FF count, and power are discussed below.
IV-C1 LUT Usage vs. Clock Frequency
IV-C2 FF Usage vs. Clock Frequency
IV-C3 Power vs. Clock Frequency
@@ IV-C Evaluation of the OBC-GEMM Module
V Performance Comparison
V-A Performance Metric Definitions

### v3

@@ III-C1 Parallel LUT
III-C2 Shared LUT
The shared LUT approach builds on the parallel LUT by sharing generated contents to reduce partial sums, as shown in Fig. 6(b). For example, LUT contents at addresses 4–7 can be derived by loading $\theta_{2}$ instead of $\theta_{2}$ via a 2-to-1 multiplexer, lowering computational cost. Directly applying this to Fig. 2(b) is challenging due to exponential adder growth. By sharing partial sums across mirrored addresses (e.g., 0–3 and 4–7), the shared LUT design reduces computations by nearly 50%.
III-C3 Split LUT
In both parallel and shared LUT techniques, adders and multiplexers grow exponentially. While shared LUTs halve computations, exponential scaling remains. To address this, the LUT is split into two parts.The first $P$ terms in the summation defined by $\tilde{Y}_{r}$ in (8) are implemented using a parallel LUT with $(2^{P-1}-1)+(P-1)$ adders and a $2^{P-1}$-to-1 multiplexer (or $2^{P-1}-1$ 2-to-1 MUXes). The remaining KPK-P terms are computed with $(2^{K-P-1}-1)+(K-P-1)$ adders and a $2^{K-P-1}$-to-1 multiplexer. One adder combines the two partial results. Thus, total adders are $(2^{P-1}+2^{K-P-1}-2)+(K-2)$ and 2-to-1 MUXes are $2^{P-1}+2^{K-P-1}-2$. Decreasing $P$ reduces first-part complexity but increases the second, and vice versa; minimizing total hardware shows the optimum at $P=K/2$. The proposed LUT architecture based on this technique is shown in Fig. 6(c). It can be observed that the chain of adders is replaced by a single adder for the first $K/2$ section, while the other term is computed using a subtractor. The remaining two partial sums are simply the twos complements of these results. A similar structure applies to the second $K/2$ section as well.
III-C4 Hybrid LUT
Although the split LUT reduces computational challenges of parallel and shared LUTs, it still grows exponentially with $K$. To address this, a hybrid LUT is proposed shown in Fig. 6(d), pairing consecutive weights and using minimal logic to generate partial sums. Here, hybrid refers to reducing large-scale parallelism via serial-like weight pairing. Adders and subtractors compute $\theta_{K-i+1}+\theta_{K-i}$ and $\theta_{K-i+1}-\theta_{K-i}$ for $i\in\{1,\dots,K-2\}$ and can be shared, while the last pair uses a conditional adder to minimize resources. Each weight pair requires only one adder or subtractor, yielding linear resource scaling. The selection lines for the 2-to-1 multiplexers to select between $\theta_{K-i+1}+\theta_{K-i}$ and $\theta_{K-i+1}-\theta_{K-i}$—can be derived from the observation that the contents at the $0^{\textnormal{th}}$ and $3^{\textnormal{rd}}$ address locations are two’s complements of each other, as are the $1^{\textnormal{st}}$ and $2^{\textnormal{nd}}$, and similarly for the second half, as illustrated in Fig. 2(a).
@@ III-D1 Hardware and Time Complexities
Table I lists theoretical hardware complexity (in terms of adders and 2-to-1 multiplexers) and time complexity (in terms of CPD) for the proposed LUT architectures, estimated using $K=p\times q$ for fair comparison. The Parallel LUT has the highest hardware cost due to exponential growth with $q$. The Shared LUT reduces adders by reusing partial sums but maintains similar CPD. The Split LUT further lowers adders at the cost of more multiplexers. The Hybrid LUT achieves the best balance, with linear adder scaling, reduced multiplexers, and low CPD due to shallow adder trees. For comparison, $q=4$ ($K=4p$), and all LUTs are implemented on FPGA with 8-bit symmetric operands, as detailed in the next subsection.
@@ III-D2 LUT and FF Usage vs Clock Frequency
All buffers are implemented using Slice RAM instead of BRAM, improving flexibility and making the architecture better suited for edge deployment while maintaining a fully LUT-based design. Fig. 7 shows the resource utilization of the proposed LUT architectures on the Xilinx ZCU106 FPGA across operating frequencies from 100 MHz to 800 MHz in 100 MHz increments. LUT usage remains stable when timing constraints are easily met but rises at higher frequencies due to tool-driven optimizations like logic duplication and register insertion, reflecting synthesis overhead rather than intrinsic design requirements. All designs show LUT usage consistent with Table I. The Parallel LUT uses the most LUTs, the Shared LUT reduces usage by reusing partial computations, the Split LUT lowers adders but increases logic via twos complementers (slightly affecting $K=32$), and the Hybrid LUT achieves the most efficient usage with optimized adders and minimal multiplexers. Minor variations arise from synthesis heuristics and routing.
Unlike LUT usage, flip-flops (FF) utilization remains largely consistent across all LUT designs, as shown in Fig. 8. This is because LUTs are primarily combinational while FFs store feature maps and pipeline data. Increasing $K$ widens the datapath, requiring more FFs for intermediate results; this increase is similar across all architectures since it depends on sequential storage rather than LUT structure.
@@ III-D3 Power Consumption vs Clock Frequency
Accurate power estimation requires realistic activity data, as post-synthesis estimates often miss switching and routing effects in complex LUTs. In Xilinx Vivado, the procedure is: (1) run functional simulation; (2) generate a .saif file with signal toggles; (3) import it into power analysis; (4) combine toggles with utilization, timing, and clock frequency. This captures actual activity and frequency behavior. Power increases with clock frequency for all LUTs due to higher switching. The slope varies with logic complexity, routing, and active elements: Parallel LUT has the steepest slope from many adders and multiplexers; Shared LUT shows low power at $K=8$ due to computation reuse; Hybrid LUT has slightly higher power at small $K$ but consistent power at $K=16$ and lowest at $K=32$ by pairing adders and minimizing multiplexers; Split LUT shows shallower slopes due to fewer adders. These trends highlight the impact of logic complexity and switching/data reuse on power.
IV Evaluation of the OBC-GEMM Module
Figure 9: LUTs/$F$s usage and power consumption of OBC-GEMM for different $(B_{1},B_{2})_{S}$ with $S\in\{A,B,AB\}$ indicates Scheme A or B or Symmetric case - AB values with hybrid and split LUT techniques at 50 MHz and 200 MHz.
For evaluation, we present the Xilinx ZCU106 results of the OBC-GEMM module with input bitwidth $B_{1}=16$ while varying the weight bitwidth $B_{2}$; the same analysis can be extended to $B_{1}=8$. Input feature maps, weights, and biases are preloaded into on-chip RAM, and the outputs are written back to the same memory. The proposed OBC-GEMM supports asymmetric quantization under Scheme A and Scheme B, as described in Section II. The selected bitwidth combinations in Fig. 9 are primarily motivated by accuracy considerations, which are analyzed in the next section. In particular, the $(16,4)_{B}$ configuration is retained to illustrate the behavior of Scheme B under aggressive weight quantization, whereas the $(16,4)$ configuration under Scheme A is omitted due to severe accuracy degradation.
Under asymmetric quantization, both proposed schemes follow $B_{1}>B_{2}$. In Scheme A, the SA unit operates for $B_{1}$ clock cycles corresponding to the input bits, while the partial products of the weights are generated using efficient hardware techniques, reducing arithmetic complexity and overall hardware resource usage. In contrast, Scheme B (Fig. 5) bit-slices the weights with bit-width $B_{2}$ through a PISO for address generation, while inputs in OBC form are used to generate the LUT data. Here, the data remain in the SA unit for $B_{2}$ clock cycles, enabling faster output, but the wider LUT data ($B_{1}$ bits) increase logic utilization and power consumption. The $(16,4)_{B}$ case illustrates this trade-off in its most extreme form: very short SA operation but significantly wider LUT paths.
After integrating the Split and Hybrid LUTs and implementing both schemes, the results are summarized in Fig. 9. To evaluate the effect of timing constraints, we vary $B_{2}$, and $K$ at two operating frequencies: 50 MHz and 200 MHz. At 50 MHz, the designs meet timing with minimal replication, while at 200 MHz additional logic and routing are required to satisfy stricter constraints. Under both conditions, Scheme A and Scheme B exhibit consistent resource–performance trade-offs, and the choice of $(B_{1},B_{2})$ directly governs efficiency. The detailed effects on LUT usage, FF count, and power consumption are discussed below.
IV-1 LUT Usage vs. Clock Frequency
IV-2 FF Usage vs. Clock Frequency
TABLE II: Description of layer types, channel counts, and activation functions in LeNet-5 and All-CNN-C
Layer LeNet-5 [21] Layer All-CNN-C [29]
Original Modified
Conv1 $5\times 5$ conv, 6 ch, tanh $5\times 5$ conv, 6 ch, ReLU Conv1 $3\times 3$ conv, 96 ch, ReLU
Pool1 $2\times 2$ avg pool, stride 2, tanh $3\times 3$ conv, stride 2, ReLU Conv2 $3\times 3$ conv, 96 ch, ReLU
Conv2 $5\times 5$ conv, 16 ch, tanh $5\times 5$ conv, 16 ch, ReLU Down1 $3\times 3$ conv, 96 ch, stride 2, ReLU
Pool2 $2\times 2$ avg pool, stride 2, tanh $3\times 3$ conv, stride 2, ReLU Conv3 $3\times 3$ conv, 192 ch, ReLU
GAP global avg pool on 16 ch Conv4 $3\times 3$ conv, 192 ch, ReLU
FC1 dense $400\!\to\!120$, tanh dense $16\!\to\!32$, ReLU Down2 $3\times 3$ conv, 192 ch, stride 2, ReLU
FC2 dense $120\!\to\!84$, tanh dense $32\!\to\!10$, softmax Conv5 $3\times 3$ conv, 192 ch, ReLU; $1\times 1$ conv, 192 ch, ReLU
FC3 dense $84\!\to\!10$, softmax Global $1\times 1$ conv, 10 ch, ReLU; global avg pool; softmax
@@ IV Evaluation of the OBC-GEMM Module
  • The All-CNN-C column summarizes the standard CIFAR-style architecture by stage, where downsampling is performed using stride-2 convolution instead of pooling.
IV-3 Power vs. Clock Frequency
@@
V Models Consideration and Accuracy Results
V-A Models Consideration
To comprehensively evaluate the applicability of the proposed OBC-GEMM framework, we consider both a classical LeNet-5 [21] and a modern All-CNN-C [29] thereby demonstrating effectiveness across shallow and deeper convolution-dominant workloads. However, the classical LeNet-5 does require lightweight architectural refinements [9], as summarized in Table II. These modifications are primarily motivated by hardware regularity, arithmetic intensity, and fixed-point robustness. Specifically:
  • Replace tanh with ReLU to avoid saturation and simplify quantized implementation.
  • Replace $2\times 2$ average pooling with $3\times 3$ stride-2 convolutions to unify computation and improve data reuse.
  • Replace the two large fully connected (FC) layers with Global Average Pooling (GAP) and a 32-unit FC layer to reduce parameters and memory traffic.
  • Insert one-side zero padding before each strided convolution to maintain spatial consistency and simplify buffering.
These modifications have minimal impact on throughput. Stride-2 convolutions increase total MACs per inference but do not lower achievable MACs per cycle. ReLU and zero-padding add negligible compute cost, while GAP reduces parameters and memory traffic without altering the core computational structure. Thus, the main benefits are improved accuracy robustness and memory efficiency rather than increased compute throughput.
We further examine All-CNN-C architecture that aligns naturally with the proposed OBC-GEMM core as listed in Table II. This architecture eliminates pooling and dense layers entirely, relying on repeated $3\times 3$ convolutions and stride-2 downsampling. From a hardware perspective, its key characteristics can be summarized as follows:
  • Downsampling is performed exclusively using stride-2 convolutions, ensuring that all layers map uniformly to GEMM operations.
  • Small $3\times 3$ kernels are used throughout, increasing nonlinearity depth while maintaining regular memory access patterns.
  • FC layers are replaced with $1\times 1$ convolutions followed by GAP and softmax, significantly reducing parameter storage.
  • ReLU is used consistently, improving fixed-point compatibility and eliminating saturation effects.
  • The repetitive convolutional structure simplifies tiling, buffering, and control logic on FPGA.
As with the modified LeNet-5, these architectural choices primarily influence arithmetic intensity and memory behavior rather than peak MAC/cycle capability. By removing heterogeneous operators such as pooling and large FC layers, All-CNN-C provides a structurally uniform workload that cleanly exercises the convolution-centric datapath of the proposed design.
V-B Accuracy Comparison Under Quantization
Prior to the FPGA implementation, we first assessed the impact of quantization-aware training (QAT) by comparing the original and modified LeNet-5 on MNIST with padded $32\times 32$ inputs. Both models were trained using fixed-point arithmetic, with input bitwidth $B_{1}\in\{16,8\}$ and varying weight bitwidth $B_{2}$, as shown in Fig. 10. When the input precision is high ($B_{1}=16$), the original LeNet-5 achieves $91.96\%$$92.58\%$ accuracy and shows minimal sensitivity to $B_{2}$. In contrast, the modified architecture consistently outperforms the baseline across all weight precisions, reaching $98.28\%$ at $B_{2}=4$ and $99.16\%$ at $B_{2}=16$. The largest gains occur at moderate weight precisions, indicating that ReLU activations, convolution-based downsampling, and GAP significantly improve robustness to weight quantization and enable near floating-point accuracy even at reduced bitwidth. Under reduced input precision ($B_{1}=8$), quantization effects become more pronounced. The original network drops to $83.18\%$ at $B_{2}=2$ but gradually recovers to $94.34\%$ at $B_{2}=8$. The modified model exhibits greater degradation at extreme low precision e.g., $64.32\%$ at $B_{2}=2$, yet performance rebounds rapidly, reaching $98.22\%$ at 4 bits and $98.60\%$ at 8 bits. At moderate precision, the modified architecture maintains a clear accuracy advantage, demonstrating strong representational capacity once a modest precision threshold is met.
We next evaluated All-CNN-C under the same QAT framework on CIFAR-10, using RGB $32\times 32$ inputs. For $B_{1}=16$, the network achieves nearly identical accuracy for weight bitwidths $B_{2}=4,8,12,$ and $16$, indicating limited sensitivity to weight precision when activation bitwidth is high. Increasing $B_{2}$ beyond 4 bits yields only marginal improvements. When $B_{1}=8$, performance drops to 76.50% at $B_{2}=2$, but recovers sharply to 85.51% at 4 bits and remains close to 85% for $B_{2}=6$–8 bits. This trend is consistent with the LeNet-5 results: ultra-low precision significantly degrades performance, whereas moderate precision is sufficient to stabilize both training and inference. Overall, these results confirm that All-CNN-C maintains stable behavior at $B_{2}\geq 4$ bits on CIFAR-10, supporting its suitability for the proposed convolution-centric OBC-GEMM framework. Across both datasets—MNIST for LeNet-5 and CIFAR-10 for All-CNN-C—moderate precision (4–8 bits) offers an effective balance between accuracy and hardware efficiency.
Figure 10: Accuracy of original/modified LeNet-5 and All-CNN-C under QAT with two sets of input bit-widths ($B_{1}$) and varying weight bit-widths ($B_{2}$).
VI Performance Evaluation and Comparison
VI-A Performance Metric Definitions